0, 1, 15, 19, 23, 31, 49, 55, 59, 87, 91, 97, 103, 105
Which of the given numbers are composite?

Answer by checkley75(3666)

15=3*5
49=7*7
55=5*11
87=3*29
91=7*13
105=3*5*7
these are all composite numbers.
that is if you exclude (1*x=x) as a composite number.